OpenStudy (anonymous):

Tanx+secx=1... Find the solutions of the equation that are in the interval [0,2pi)

OpenStudy (anonymous):

Use the fact that: \[\tan x=\frac{\sin x}{\cos x}\] and \[\sec x=\frac{1}{\cos x}\] We can change substitute the values on the RHS in your equation. Your ORIGINAL equation was this: \[\tan x +\sec x=1\] Your ADJUSTED equation is now this when you substitute both tanx and secx with the values I've listed: \[\frac{\sin x}{\cos x}+\frac{1}{\cos x}=1\] Now you combine the two fractions together since both fractions have the same denominator: \[\frac{\sin x+1}{\cos x}=1\] Now we can move cosx to the RHS, to give you this: \[\sin x+1=\cos x\] Next step is to put sinx and cosx together on one side and the 1 on its own to the other side: \[1=\cos x -\sin x\] \[\cos x-\sin x =1\] Now there is a method known as the Auxiliary angle method. You have to get the above equation in the form of this: \[Rcos(x+\epsilon)=1\] where: the auxiliary angle is the "e" symbol or commonly known as epsilon \[R=\sqrt{a^2+b^2}\] \[a\cos x\pm b\sin x=1\] If you need help with the auxiliary angle method, it's best to ask your teacher when you get the time in class to do so.